OpenStudy (anonymous):

What is the slope of the line whose equation is 4x – 6y = 12?

OpenStudy (anonymous):

solve for y and m is the slope in y=mx+b

OpenStudy (anonymous):

a)-4 b)-2/3 c)2/3 d) 6

OpenStudy (anonymous):

c

OpenStudy (anonymous):

or d

OpenStudy (anonymous):

c sounds legit to me but why would you think d?

OpenStudy (nurali):

y=mx+b 4x – 6y = 12 First find the value of y. 6y=4x-12 y=4x-12/6 y=4x/6-12/6 y=2x/3-2

OpenStudy (anonymous):

i got -12 for d ...... i dont know lol @lonnie455rich

OpenStudy (anonymous):

thanks @Nurali :)

OpenStudy (nurali):

Anytime.

OpenStudy (anonymous):

c or b

OpenStudy (anonymous):

lol c you should make sure you understand why though.

OpenStudy (anonymous):

c

OpenStudy (anonymous):

lol ikr @lonnie455rich :D

OpenStudy (nurali):